18. The physician orders dopamine for the client. The pharmacy sends an IV containing 200 mg in 500mL NS. The order states: Begin titration at 5mcg/kg/min and client weighs 132lbs. Calculate the initial flow rate. The client will receive __________mL/hr (Round to whole number)
Answer #1

Available dose

200mg= 500ml

2mg= 5ml

Weight= 132lb= 132/2.2= 60kg

Order

5MCG/kg/min

5MCG×60/min

300MCG/min

Hour

300×60/hour

18000MCG/hour

18mg/hour

available dose 2mg=5ml

2×9mg= 9×5ml

18mg=45ml

Answer

45ml/hour